a=Square root of 7a-6
Found 2 solutions by checkley77, nycsharkman:
Answer by checkley77(12844)

A=SQRT(7A-6)
SQUARE BOTH SIDES OF THIS EQUATION
A^2=7A-6
A^2-7X+6=0
(A-6)(A-1)=0
A-6=0
A=6 ANSWER.
A-1=0
A=1 ANSWER.

Answer by nycsharkman(136)

a=Square root of 7a-6
I guess you are trying to solve for a , right?
Square both sides:
(a)^2 = (sqrt{7a - 6})^2
a^2 = 7a - 6
Bring the right side to the left side and equate to zero.
a^2 - 7a + 6 = 0
Factor left side.
(a - 6)(a - 1) = 0
Set each factor to zero and solve for a.
a - 6 = 0
a = 6
=========
a - 1 = 0
a = 1
Final answer: a = 1 and a = 6
We have to check the answers for a because we are dealing with square roots.
CHECKING:
a = sqrt{7a - 6}
I will plug 1 for a and simplify and then plug 6 for a and simplify. I must get the same answer on BOTH sides of the equation.
Let a = 1
1 = sqrt{7(1) - 6}
1 = sqrt{7 - 6}
1 = sqrt{1}
1 = 1...IT CHECKS!
==================
Now, let a = 6 and do the same thing again.
6 = sqrt{7(6) - 6}
6 = sqrt{42 - 6}
6 = sqrt{36}
6 = 6...IT CHECKS!
So, now I know for sure that a = 1 and a = 6.
Is this clear?